OpenStudy (anonymous):

inverse f(x)= (2x+1) / (3x+2)

OpenStudy (anonymous):

\[f(x)=(2x+1)\div (3x+2) \] right?

OpenStudy (anonymous):

yeah the inverse

OpenStudy (anonymous):

of that please

OpenStudy (anonymous):

If so it would be (3x+2)(2)-( 2x-1)(3) / \[(3x+2)^{2}\]

OpenStudy (helder_edwin):

do you mean the inverse function for \[ \Large f(x)=\frac{2x+1}{3x+2} \]

OpenStudy (anonymous):

yes ^

OpenStudy (anonymous):

yes but thats not the anserw but the first step I did

OpenStudy (helder_edwin):

ok then you have to solve for x the equation \[ \Large y=f(x)=\frac{2x+1}{3x+2} \] tell me what you get

OpenStudy (anonymous):

would the anserw be 6 over \[(3x+2)^{2}\]

OpenStudy (helder_edwin):

No @hooverst

OpenStudy (helder_edwin):

@sabriiina. are you working on this?

OpenStudy (anonymous):

how would u do it? I probably missed a step

OpenStudy (anonymous):

i know what the answer is

OpenStudy (anonymous):

i don't know how to get it

OpenStudy (anonymous):

im stuck at the step y= 2y+1 -2x ------------- 3x

OpenStudy (helder_edwin):

ok let me try \[ \Large y=\frac{2x+1}{3x+2} \] \[ \Large y(3x+2)=2x+1 \] \[ \Large 3xy+2y=2x+1 \] \[ \Large 3xy-2x=1-2y \] \[ \Large x(3y-2)=1-2y \] \[ \Large x=\frac{1-2y}{3y-2} \] then \[ \Large f^{-1}(y)=\frac{1-2y}{3y-2} \]

OpenStudy (helder_edwin):

do you get this?

OpenStudy (valpey):

\[y=\frac{2x+1}{3x+2}\] \[({3x+2})y={2x+1}\] \[{3xy+2y}={2x+1}\] \[3xy-2x=1-2y\] \[x(3y-2)=1-2y\] \[x=\frac{1-2y}{3y-2}\]

OpenStudy (anonymous):

thank you. i get it now :)